OpenStudy (anonymous):

how to find the f(x) of the derivative of a function? f(x) = 5x-2/2x+7

OpenStudy (anonymous):

@ganeshie8

OpenStudy (luigi0210):

You could use quotient rule

OpenStudy (luigi0210):

\[\LARGE \frac{d}{dx} \frac{f(x)}{g(x)}=\frac{g(x)*f'(x)-g'(x)*f(x)}{(g(x))^2}\]